# Refinements of the Blanco-Koldobsky-Turnšek Theorem

### v1

@@ 2. Main Results
Let $F$ be a $k$-face of $B_{\mathbb{X}}.$ Consider the subspace $\mathbb{V}={\rm span}\{u-v:u,v\in F\}.$ Clearly, ${\rm dim}\,\mathbb{V}=k.$ Then $F^{\prime}={\rm span}\{u-x:u\in F\}$ is a convex polyhedron contained in $\mathbb{V}$ of dimension $k.$ Since $x$ is an extreme point of $F,$ it follows that 00 is an extreme point of $F^{\prime}.$ Let $G^{\prime}$ be a $(k-1)$-face of $F^{\prime}$ containing $0$ Now let us consider the set $G={\rm span}\{w+x:w\in G^{\prime}\}.$ Then clearly $G\subset F$ is a convex polyhedron of dimension $k-1$ containing $x.$ Now we show that $G$ is a face of $F.$ Let $y\in G$ be such that $y=(1-t)y_{1}+ty_{2}$ for $y_{1},y_{2}\in F$ and for $t\in(0,1).$ Now $y\in G\implies y-x\in G^{\prime}$ and $y_{1},y_{2}\in F\implies y_{1}-x,y_{2}-x\in F^{\prime}.$ This shows that
@@
Since $G^{\prime}$ is a face of $F^{\prime},$ it follows that $y_{1}-x,y_{2}-x\in G^{\prime}$ and so $y_{1},y_{2}\in G.$ Thus, $G$ is a face of $F$ of dimension $k-1$ containing $x$ and consequently, $G$ is a $(k-1)$-face of $B_{\mathbb{X}}$ such that $x\in G\subset F.$
@@
Let $x\in{\rm Ext}\,F.$ From Proposition 2.25, it follows that there exists a $(n-k)$-face $G$ of $B_{\mathbb{X}}$ such that $x\in G\subset F.$ Let $y\in{\rm Int_{r}}\,G$ and so $y\in{\rm k\text{-}Sm}\,\mathbb{X}\cap F$ such that $J(y)\subset J(x).$ Consider the sequence $\{x_{n}\},$ where $x_{n}=(1-\frac{1}{n})x+\frac{1}{n}y$ for all $n\in\mathbb{N}.$ Clearly, $x_{n}\in F$ for all $n\in\mathbb{N}$ and $x_{n}\longrightarrow x,$ as $n\longrightarrow\infty.$ Now we show that $J(x_{n})=J(y)$ for all $n\in\mathbb{N}.$ Let $f\in J(y),$ then $f\in J(x).$ Hence
@@
$Z$$X$$Y$$(1,0,0)$00$(0,0,1)$$(0,1,0)$
[…]

### v2

@@ 2. Main Results
Let $F$ be a $k$-face of $B_{\mathbb{X}}.$ Consider the subspace $\mathbb{V}={\rm span}\{u-v:u,v\in F\}.$ Clearly, ${\rm dim}\,\mathbb{V}=k.$ Then $F^{\prime}={\rm span}\{u-x:u\in F\}$ is a convex polyhedron contained in $\mathbb{V}$ of dimension $k.$ Since $x$ is an extreme point of $F,$ it follows that 0 is an extreme point of $F^{\prime}.$ Let $G^{\prime}$ be a (k1)(k-1)-face of $F^{\prime}$ containing $0.$ Now let us consider the set $G={\rm span}\{w+x:w\in G^{\prime}\}.$ Then clearly $G\subset F$ is a convex polyhedron of dimension k1k-1 containing $x.$ Now we show that $G$ is a face of $F.$ Let $y\in G$ be such that $y=(1-t)y_{1}+ty_{2}$ for $y_{1},y_{2}\in F$ and for $t\in(0,1).$ Now $y\in G\implies y-x\in G^{\prime}$ and $y_{1},y_{2}\in F\implies y_{1}-x,y_{2}-x\in F^{\prime}.$ This shows that
@@
Since $G^{\prime}$ is a face of $F^{\prime},$ it follows that $y_{1}-x,y_{2}-x\in G^{\prime}$ and so $y_{1},y_{2}\in G.$ Thus, $G$ is a face of $F$ of dimension k1k-1 containing $x$ and consequently, $G$ is a (k1)(k-1)-face of $B_{\mathbb{X}}$ such that $x\in G\subset F.$
@@
Let $x\in{\rm Ext}\,F.$ From Proposition 2.25, it follows that there exists a (nk)(n-k)-face $G$ of $B_{\mathbb{X}}$ such that $x\in G\subset F.$ Let $y\in{\rm Int_{r}}\,G$ and so $y\in{\rm k\text{-}Sm}\,\mathbb{X}\cap F$ such that $J(y)\subset J(x).$ Consider the sequence $\{x_{n}\},$ where $x_{n}=(1-\frac{1}{n})x+\frac{1}{n}y$ for all $n\in\mathbb{N}.$ Clearly, $x_{n}\in F$ for all $n\in\mathbb{N}$ and $x_{n}\longrightarrow x,$ as $n\longrightarrow\infty.$ Now we show that $J(x_{n})=J(y)$ for all $n\in\mathbb{N}.$ Let $f\in J(y),$ then $f\in J(x).$ Hence
@@
ZZXXYY(1,0,0)(1,0,0)0(0,0,1)(0,0,1)$(0,1,0)$
This unit sphere is a surface of revolution obtained by rotating a square with vertices $\{\pm(1,0,0),\pm(0,0,1)\}$ around the $X$-axis. Here $2\text{-Sm~}\mathbb{X}=\{(0,b,c):b,c\in\mathbb{R},(0,b,c)\neq(0,0,0)\}$ and $3\text{-Sm~}\mathbb{X}=\{(a,0,0):a\in\mathbb{R}\setminus\{0\}\}.$ Since $\dim{\rm span}(2\text{-Sm~}\mathbb{X})=2$ and $\dim{\rm span}(3\text{-Sm~}\mathbb{X})=1,$ it follows from [21, Prop. 2.1] that $2\text{-Sm~}\mathbb{X}$ and $3\text{-Sm~}\mathbb{X}$ are not $\mathcal{K}$-sets.
@@
    $T$ preserves Birkhoff-James orthogonality at each point of $U,$ where $U$ is a dense subset of $S_{\mathbb{X}}.$
@@
Conflict of interest
The authors have no relevant financial or non-financial interests to disclose. The authors have no competing interests to declare that are relevant to the content of this article.